OpenStudy (anonymous):

simplify: a) 12x3^4/6x3^2 b)3^4/3 c) 8/2^4 d) (-3w)^2 e) (-2t^6)^3

OpenStudy (anonymous):

ok, looks like a through c just have to deal with common factors

OpenStudy (anonymous):

ok, looks like a through c just have to deal with common factors

OpenStudy (anonymous):

for example in (a): \[12*3^4 / 6*3^2\] \[3^4 = 3*3*3*3 = 3^2 * 3^2\]

OpenStudy (anonymous):

so \[12*3^4 = 12*3^2*3^2 \] then \[(12 * 3^2 * 3^2) / ( 6 * 3^2) = (12 * 3^2)/ 6 = (12 / 6) * (3 ^2) = 2 * 3^2\]

OpenStudy (anonymous):

so the final = is

OpenStudy (anonymous):

2*3^2